OpenStudy (mrchair):

Solve for x. -3x - 8 + 6x = 7 x = - 7 /5 x = - 1/ 3 x = 5 x = 45

OpenStudy (mrchair):

those are fractions, ill fix

OpenStudy (kirbykirby):

\[ -3x - 8 + 6x = 7 \\ -3x + 6x - 8 = 7\\ 3x-8=7\\ 3x=7+8\\ 3x=15\\ x=\frac{15}{3}\\ x=5\]

OpenStudy (mrchair):

wow thanks, i missed a step

OpenStudy (kirbykirby):

=]

OpenStudy (mrchair):

ill have a lil more if u wanna help